OpenStudy (anonymous):

An artillery shell is fired at an angle of 62.7 above the horizontal ground with an initial speed of 1520 m/s. The acceleration of gravity is 9.8 m/s2 . Find the total time of flight of the shell, neglecting air resistance. Answer in units of min

OpenStudy (anonymous):

vertical is 1520sin62.7 = 1350.69819 horizontal is 1520cos62.7 = 697.14732 s=ut +1/2 at^2 therefore 1350.698t +4.9t^2 t=277.77 seconds... and then divide 697.1432/277.77 = part b Can someone tell me if this is the correct answer, or if not what I am doing wrong.

OpenStudy (anonymous):

oh yeah this is part B lol... Find its horizontal range, neglecting air resis- tance. Answer in units of km

OpenStudy (anonymous):

It appears you know what you're doing. Check this wikipedia article for an equation reference: http://en.wikipedia.org/wiki/Projectile_motion

OpenStudy (anonymous):

I always make small eroors in my calculations and I end up with the wrong answer. Did you by chance check the work to see if it is correct. I didnt want to enter it into my HW assignment until I know for sure it is correct.

OpenStudy (anonymous):

errors*

OpenStudy (anonymous):

Double checked and verified. Great work.

OpenStudy (anonymous):

It was wrong :(